On january 1, canglon, inc., issues 19%, 5-year binds with a face value if $150,000 when the effective rate is 12%. Interest is to be paid semiannually.

Prepare calculations to prove that the selling price of the bonds is $138,959.90.

What is the present value if principal?
What is the present value if interest?
Answer #1
Bond characterstics Amount
1-a) Principal 150,000
interest                7,500
Market interest rate 6%
periods to maturity 10
issue price 138,959
Calculation of bond issue price
Where
i= 6.00%
t= 10
principal * PV of $1 at 6% for 10 yrs =
150,000 * 0.55839        = 83758.5
interest * PV of ordinary annuity at 6%=
7500 * 7.36009 = 55200.4
bond issue price 138958.9